Cathy and Tom's Specialty Ice Cream Company operates a small production facility for the local community. The facility has the capacity to make 16,400 gallons of the single flavor, GUI Chewy, annually. The plant has only two customers, Chuck's Gas & Go and Marcee's Drive & Chew DriveThru. Annual orders for Chuck's total 9,500 gallons and annual orders for Marcee's total 3,800 gallons. Variable manufacturing costs are $0.90 per gallon, and annual fixed manufacturing costs are $24,500. Required: What cost per gallon should the cost system report? (Round your intermediate calculations and final answers to 2 decimal places.)

Cost per gallon at capacity

cost per gallon at demand
Answer #1

Ans:

Cost per gallon = Total cost / Number of gallons

Cost per gallon At capacity:

The facility has the capacity to make 16,400 gallons.

Total cost i.e(Fixed+ variable cost): 24,500 + 16,400 gallons*$0.9 = $39,260
Cost per gallon = $39,260/16,400 = $2.39 per gallon

cost per gallon At demand:

you make only 9,500 + 3,800 = 13,300 gallons.

Total costs = 24,500 + 13,300*$0.9 = $36,470
Cost per gallon = $36,470/13,300 = $2.74

Note:

1)Fixed cost never changes with respect to production

2)variable cost per unit constant